# Reliable-loc: Robust sequential LiDAR global localization in large-scale street scenes based on verifiable cues

### v1

@@ Abstract
Wearable laser scanning (WLS) system has the advantages of flexibility and portability. It can be used for determining the user’s path within a prior map, which is a huge demand for applications in pedestrian navigation, collaborative mapping, augmented reality, and emergency rescue. However, existing LiDAR-based global localization methods suffer from insufficient robustness, especially in complex large-scale outdoor scenes with insufficient features and incomplete coverage of the prior map. To address such challenges, we propose LiDAR-based reliable global localization (Reliable-loc) exploiting the verifiable cues in the sequential LiDAR data. First, we propose a Monte Carlo Localization (MCL) based on spatially verifiable cues, utilizing the rich information embedded in local features to adjust the particles’ weights hence avoiding the particles converging to erroneous regions. Second, we propose a localization status monitoring mechanism guided by the sequential pose uncertainties and adaptively switching the localization mode using the temporal verifiable cues to avoid the crash of the localization system. To validate the proposed Reliable-loc, comprehensive experiments have been conducted on a large-scale heterogeneous point cloud dataset consisting of high-precision vehicle-mounted mobile laser scanning (MLS) point clouds and helmet-mounted WLS point clouds, which cover various street scenes with a length of over 20km. The experimental results indicate that Reliable-loc exhibits high robustness, accuracy, and efficiency in large-scale, complex street scenes, with a position accuracy of 1.66m, yaw accuracy of 3.09 degrees, and achieves real-time performance. For the code and detailed experimental results, please refer to https://github.com/zouxianghong/Reliable-loc.
@@ 1 Introduction
In real applications, the WLS-based localization system needs to be operated in various scenes, posing many challenges to the reliability of the global localization algorithm. First, the localization system faces challenges from feature-insufficient environments. Take the typical urban street scene in southern China as an example: the road is wide and full of viaducts, and many road sections are only flanked by street trees and flat facades, lacking sufficient features (Li et al., 2023; Lin et al., 2023). Second, the prior map may not fully cover the area, and blank data holes exist. For example, the prior point cloud map is usually collected along roads by the vehicle-mounted MLS systems and can not fully cover roadside areas due to occlusion and accessibility (Serna and Marcotegui, 2013; Mi et al., 2021). In recent years, some LiDAR-based sequential global localization methods have integrated place recognition techniques into MCL and are applicable to large-scale outdoor scenes (Yin et al., 2023a). However, existing methods still have limitations facing the above challenges. On the one hand, most of them solely rely on the global feature (the aggregation of local features) extracted from local maps to construct the observation model in MCL. In feature-insufficient scenes where global features lack descriptive adequacy, particles in MCL are prone to converging towards erroneous regions, ultimately resulting in the localization system crashing. While many place recognition methods extract local features simultaneously to aggregate the global feature, the rich information embedded in local features is ignored by the downstream MCL task. On the other hand, existing methods only rely on point cloud registration for localization once MCL converges, thus leading to localization failure due to continuous unreliable pose estimation in the scenes with insufficient features and incomplete map coverage.
@@ 3 Problem Definition and Preliminary
Let the prior point cloud map be $M=\left\{M_{i}\right\}$, where $M$ is a set consisting of point cloud submaps in the map. The sequence of point cloud submaps $Z_{1:t}$ are acquired by the WLS system up to time $t$, and its motions $U_{1:t}$ are up to time $t$. The estimation of the system’s poses $\hat{X}_{1:t}$, in the global localization task can be formulated in terms of Bayes’ law as follows:
@@ 3.2 Spectral matching
As demonstrated in Leordeanu and Hebert (2005), the approximated optimal solution $V^{*}$ can be calculated as the principal eigenvector of $A$ by the Raleigh’s ratio theorem (Parlett, 1998). Then the inter-cluster score corresponding to the optimal solution $S^{*}$ can be calculated by:
@@ 4.1 Spatial verification-based Monte Carlo Localization
$p_{G}(Z_{t}|X_{t},M)\propto F_{G}(Z_{t},M_{i}),$ (9)
Where $F_{G}$ extracts the global features of $Z_{t}$ and $M_{i}$ and calculates their similarity, describing the overall similarity of the scenes where the system and the particles are located. In general, the information embedded in the global feature is very limited and insufficient to ensure that particles converge in the correct direction, especially in feature-insufficient scenes.
@@
$p_{SM}(Z_{t}|X_{t},M)\propto\frac{s_{k}^{*}}{s_{max}^{*}},$ (10)
where $S_{max}^{*}$ is the maximum of the inter-cluster scores for all particle clusters. It should be noted that all particles in $C_{k}$ use the same observation model to adjust their weights. Meanwhile, the 3-DOF pose of the system $X_{t}^{*}=\{x_{t}^{*},y_{t}^{*},yaw_{t}^{*}\}$ can be estimated based on the correspondences corresponding to $S^{*}$ using SVD (Singular Value Decomposition) (Golub and Reinsch, 1971).
@@ 5.1 Experiment data
In this paper, the proposed method’s effectiveness is verified on a heterogeneous point cloud dataset collected by a vehicle-mounted MLS system and a helmet-mounted WLS system in Wuhan, China. High-precision vehicle-mounted MLS point clouds serve as the prior map and helmet-mounted WLS point clouds are used for localization. The dataset contains six data: CS college, Info campus, Zhongshan park, Jiefang road, Yanjiang road 1, Yanjiang road 2, as shown in Fig.7. The first two data are collected on a college campus with abundant vegetation, where roadside buildings are often obscured by trees. The road of CS college is relatively open, whereas the Info campus exhibits a more unstructured environment, with numerous blank data holes present in the prior map. The last four data are collected on urban roads with regularly arranged roadside trees, lots of high-rise buildings and viaducts in the scene, and lots of dynamic objects on the road, with variations in some of the scenes. Some of the road sections in the last four data have insufficient features, posing a significant challenge for localization. The details of creating the dataset are described in our previous work, PatchAugNet (Zou et al., 2023). The difference lies in the fact that the helmet point cloud submaps are generated every 0.5m along the trajectory and the MLS point cloud submaps are derived by utilizing 5x5m regular grids. Additionally, we introduce noise into the helmet trajectory by referencing the official code of Overlap-loc (Chen et al., 2020). The comprehensive details of the experimental data are presented in Table 1.
@@
width = colspec = Q[140]Q[240]Q[90]Q[100]Q[100]Q[140]Q[140], cells = c, cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=20.15, cell16 = r=2, cell17 = r=2, cell52 = r=4, hline1,3,9 = -0.08em, Datasets & Scene description Difficult level Acquisition time Query trajectory length (km) Number of map submaps
Query Map
CS college University campus road. Easy 2022-07-29 2021-11-20 0.9 2493
Info campus University campus road, partial map isincomplete. Hard 2021-11-12 2021-11-20 2.6 5878
Zhongshan park Urban road, the features of some road sections are insufficient. Medium 2022-08-23 2019-12-11 4.3 15186
Jiefang road Medium 2022-08-23 2020-03-20 3.2 12202
Yanjiang road 1 Hard 2022-08-23 2020-03-22 1.7 13772
Yanjiang road 2 Medium 2022-08-23 2020-03-22 2.9 13772
@@ 5.2 Experiment setting
Implementary details. We use the official model supplied by PatchAugNet (Zou et al., 2023) to extract global and local features. The spectral matching method described in Subsection 3.2 utilizes the official code of SGV (Vidanapathirana et al., 2023), renowned for its excellent parallelism capabilities. Furthermore, the local pose estimation in Subsection 4.2 is achieved based on the official code of Teaser++ (Yang et al., 2020), renowned for its robust performance. The number of particles initialized for MCL is 5000, while after convergence, it reduces to 400. In particle clustering, the distance threshold $d_{cluster}$ is set to 5m, and $K_{thr}$ is set to 40. The parameters of the Gaussian kernel function in Eq. (11) are set to 30m, 30m, and 60 degrees, respectively. For localization status monitoring, the thresholds $\sigma_{x_{thr}}$, $\sigma_{y_{thr}}$, and $\sigma_{yaw_{thr}}$ are set to 30m, 30m, and 15 degrees, respectively. $\lambda_{thr}$ is assigned a value of 1e-4 on Yanjiang road 1, and 5e-4 for the remaining datasets. All experimental methods are implemented in Python, and all experiments are conducted on an Intel® Core i7-13700KF CPU and an Nvidia® GeForce RTX 3080Ti GPU.
Evaluation metrics. To evaluate the proposed method’s effectiveness, we compare the estimated trajectories with the ground truth trajectories and then calculate the mean error and root mean square error (RMSE) of position and yaw angle. It should be noted that only results after MCL converges are considered. Additionally, we count the proportion of position errors less than 2, 5, 10, 15, 20, 25, and 30 meters and yaw angle errors less than 2, 5, 10, 15, 20, 25, and 30 degrees, i.e., the localization success rate under different position and yaw error thresholds. For convenience, we denote $Rate@XmYdeg$ as the proportion of position errors less than $X$m and yaw errors less than $Y$ degrees. Similarly, $Rate@Xm$ denotes the proportion of position errors less than $X$m.
@@ 5.3 Quantitative evaluation
Compared with PF-loc, $Rate@2m5deg$, position accuracy, and yaw accuracy of PF-SGV-loc are 3.44 percent, 32.46 m, and 6.45 degrees higher, respectively. Similarly, $Rate@2m5deg$, position accuracy, and yaw accuracy of PF-SGV2-loc are 18.11 percent, 1.50 meters, and 2.92 degrees higher than that of PF-SGV-loc, respectively. The reason for performance improvement is that the spatial verification-based observation model can effectively utilize the rich information embedded in local features, which is an effective complement to the observation model based on global features. Specifically, in feature-rich scenes, higher localization accuracy can be achieved with the spatial-verification-based observation model, providing a better initial pose for subsequent localization. In feature-insufficient scenes, higher localization robustness can be achieved by avoiding the particles converging to the incorrect region.
PF-SGV2-loc presents a high $Rate@30m$ on all data, exhibiting excellent robustness. However, the corresponding position and yaw accuracy are likely as low as 35.15m and 11.65 degrees, respectively, exhibiting poor accuracy. Reg-loc achieves position and yaw accuracies of 1.01m and 2.51 degrees respectively on the simple data (CS college), exhibiting excellent accuracy. While it presents a low $Rate@30m$ in other data with insufficient features, exhibiting poor robustness. Reliable-loc outperforms both PF-SGV2-loc and Reg-loc on all data with a total position accuracy and yaw accuracy of 1.66m and 3.09 degrees, respectively, and it also presents a high $Rate@30m$ on all data, exhibiting excellent robustness and accuracy. The experimental results show that MCL is robust but not accurate, and localization based on point cloud registration is effective in feature-rich scenes but prone to failure in scenes with insufficient features and incomplete map coverage. However, Combining the two methods significantly improves the localization robustness and accuracy. The reasons are as follows: 1) the particles in MCL possess certain exploratory capabilities, allowing them to converge to the correct region even when the initial pose is inaccurate.; 2) After MCL converges, local pose estimation can achieve high localization accuracy in scenes with sufficient distinguishable features and provide an accurate initial pose for subsequent localization; 3) Reliable-loc can improve localization performance by adaptively switching the localization mode based on pose uncertainty from correspondences and odometry. It uses the $Reg$ localization mode when the initial pose is accurate and there are abundant features in the scene while using the $PF$ localization mode in other situations.
@@
width = colspec = Q[90]Q[140]Q[80]Q[102]Q[115]Q[102]Q[110]Q[120]Q[102], cell21 = r=3, cell51 = r=3, cell81 = r=3, cell111 = r=3, cell141 = r=3, hline1-2,5,8,11,14,17 = -0.08em, Methods & Metrics CS college Info campus Zhongshan park Jiefang road Yanjiang road 1 Yanjiang road 2 Total
PF-loc Rate@2m5deg (%)↑ 4.76 0.39 5.14 1.50 1.50 4.29 3.00
Position Error (m)↓ 8.96 ± 10.41 241.05 ± 289.59 8.28 ± 14.43 7.77 ± 10.49 7.65 ± 8.91 12.12 ± 17.06 42.23 ± 110.33
Yaw Error (deg)↓ 7.98 ± 9.49 53.26 ± 67.94 8.25 ± 10.73 7.90 ± 10.48 7.64 ± 9.24 9.90 ± 12.26 14.86 ± 27.59
PF-SGV-loc Rate@2m5deg (%)↑ 0.81 2.33 7.00 9.24 6.58 7.96 6.44
Position Error (m)↓ 7.16 ± 7.95 23.49 ± 76.67 6.34 ± 12.92 5.32 ± 6.39 5.38 ± 7.27 12.66 ± 20.43 9.77 ± 31.75
Yaw Error (deg)↓ 6.68 ± 7.96 10.91 ± 22.78 7.92 ± 10.31 7.31 ± 9.34 6.96 ± 8.80 9.83 ± 13.29 8.41 ± 13.06
PF-SGV2-loc Rate@2m5deg (%)↑ 11.25 17.77 23.11 34.75 27.00 24.45 24.55
Position Error (m)↓ 3.38 ± 3.73 35.15 ± 102.23 3.22 ± 10.71 2.57 ± 3.16 3.31 ± 4.11 4.43 ± 6.73 8.27 ± 40.77
Yaw Error (deg)↓ 4.17 ± 5.20 11.65 ± 28.36 4.38 ± 5.58 3.80 ± 4.84 4.53 ± 5.97 4.91 ± 6.64 5.49 ± 12.33
Reg-loc Rate@2m5deg (%)↑ 73.78 6.64 55.37 42.95 2.71 2.00 29.92
Position Error (m)↓ 1.01 ± 1.83 164.37 ± 201.16 113.08 ± 227.61 458.15 ± 773.78 762.37 ± 938.22 1356.16 ± 1579.88 489.74 ± 840.14
Yaw Error (deg)↓ 2.51 ± 4.37 32.22 ± 39.78 9.82 ± 18.26 34.25 ± 54.00 65.78 ± 75.52 58.02 ± 63.06 33.36 ± 48.26
@@
Rate@2m5deg (%)↑ 77.55 52.02 80.39 72.37 35.91 68.87 66.71
Position Error (m)↓ 0.73 ± 1.10 2.41 ± 3.31 1.13 ± 10.29 1.92 ± 5.34 2.39 ± 3.09 1.33 ± 2.59 1.66 ± 6.29
Yaw Error (deg)↓ 2.09 ± 3.55 4.02 ± 6.12 2.44 ± 4.57 2.64 ± 4.42 4.45 ± 6.02 3.22 ± 5.23 3.09 ± 5.08 Note: the rightmost column ’Total’ is the overall localization performance of each method on all experimental data.
Figure 8: Localization success rate curves of each method on experimental data. (a), (b), (c), (g), (h), (i) are the $Rate@Xm$ curves, and (d), (e), (f), (j), (k), (l) are the $Rate@Ydeg$ curves.

### v2

@@ Abstract
Wearable laser scanning (WLS) system has the advantages of flexibility and portability. It can be used for determining the user’s path within a prior map, which is a huge demand for applications in pedestrian navigation, collaborative mapping, augmented reality, and emergency rescue. However, existing LiDAR-based global localization methods suffer from insufficient robustness, especially in complex large-scale outdoor scenes with insufficient features and incomplete coverage of the prior map. To address such challenges, we propose LiDAR-based reliable global localization (Reliable-loc) exploiting the verifiable cues in the sequential LiDAR data. First, we propose a Monte Carlo Localization (MCL) based on spatially verifiable cues, utilizing the rich information embedded in local features to adjust the particles’ weights hence avoiding the particles converging to erroneous regions. Second, we propose a localization status monitoring mechanism guided by the sequential pose uncertainties and adaptively switching the localization mode using the temporal verifiable cues to avoid the crash of the localization system. To validate the proposed Reliable-loc, comprehensive experiments have been conducted on a large-scale heterogeneous point cloud dataset consisting of high-precision vehicle-mounted mobile laser scanning (MLS) point clouds and helmet-mounted WLS point clouds, which cover various street scenes with a length of over 30 km. The experimental results indicate that Reliable-loc exhibits high robustness, accuracy, and efficiency in large-scale, complex street scenes, with a position accuracy of ±2.91 m, yaw accuracy of ±3.74 degrees, and achieves real-time performance. For the code and detailed experimental results, please refer to https://github.com/zouxianghong/Reliable-loc.
@@ 1 Introduction
In real applications, the WLS-based localization system needs to be operated in various scenes, posing many challenges to the reliability of the global localization algorithm. First, the localization system faces challenges from feature-insufficient environments. Take the typical urban street scene in southern China as an example: the road is wide and full of viaducts, and many road sections are only flanked by street trees and flat facades, lacking sufficient features (Li et al., 2023; Han et al., 2024). Second, the prior map may not fully cover the area, and blank data holes exist. For example, the prior point cloud map is usually collected along roads by the vehicle-mounted MLS systems and can not fully cover roadside areas due to occlusion and accessibility (Serna and Marcotegui, 2013; Mi et al., 2021). In recent years, some LiDAR-based sequential global localization methods have integrated place recognition techniques into MCL and are applicable to large-scale outdoor scenes (Yin et al., 2023a). However, existing methods still have limitations facing the above challenges. On the one hand, most of them solely rely on the global feature (the aggregation of local features) extracted from local maps to construct the observation model in MCL. In feature-insufficient scenes where global features lack descriptive adequacy, particles in MCL are prone to converging towards erroneous regions, ultimately resulting in the localization system crashing. While many place recognition methods extract local features simultaneously to aggregate the global feature, the rich information embedded in local features is ignored by the downstream MCL task. On the other hand, existing methods only rely on point cloud registration for localization once MCL converges, thus leading to localization failure due to continuous unreliable pose estimation in the scenes with insufficient features and incomplete map coverage.
@@ 3 Problem Definition and Preliminary
Let the prior point cloud map be $M=\left\{M_{i}\right\}$, where $M$ is a set consisting of point cloud submaps in the map. The sequence of point cloud submaps $Z_{1:t}$ are acquired by the WLS system up to time $t$, and its motions $U_{1:t}$ are up to time $t$. The estimation of the system’s poses $\hat{X}_{1:t}$, in the global localization task, can be formulated in terms of Bayes’ law as follows:
@@ 3.2 Spectral matching
As demonstrated by Leordeanu and Hebert (2005), the approximated optimal solution $V^{*}$ can be calculated as the principal eigenvector of $A$ by the Raleigh’s ratio theorem (Parlett, 1998). Then the inter-cluster score corresponding to the optimal solution $S^{*}$ can be calculated by:
@@ 4.1 Spatial verification-based Monte Carlo Localization
$p_{G}(Z_{t}|X_{t},M)\propto F_{G}(Z_{t},M_{i})=\frac{\cos(f_{Z_{t}},f_{M_{i}})+1}{2},$ (9)
Where $F_{G}$ extracts the global features of $Z_{t}$ and $M_{i}$ and calculates their similarity, describing the overall similarity of the scenes where the system and the particles are located. $f_{Z_{t}}$ and $f_{M_{i}}$ are the global features of $Z_{t}$ and $M_{i}$, respectively. In general, the information embedded in the global feature is very limited and insufficient to ensure that particles converge in the correct direction, especially in feature-insufficient scenes.
@@
$p_{SM}(Z_{t}|X_{t},M)\propto\frac{s_{k}^{*}}{s_{max}^{*}},~{}S_{max}^{*}=\max_{m}S_{m}^{*},$ (10)
where $S_{max}^{*}$ is the maximum of the inter-cluster scores for all particle clusters, $m$ is the index of the particle cluter. It should be noted that all particles in $C_{k}$ use the same observation model to adjust their weights. Meanwhile, the 3-DOF pose of the system $X_{t}^{*}=\{x_{t}^{*},y_{t}^{*},yaw_{t}^{*}\}$ can be estimated based on the correspondences corresponding to $S^{*}$ using SVD (Singular Value Decomposition) (Golub and Reinsch, 1971).
@@ 5.1 Experiment data
In this paper, the proposed method’s effectiveness is verified on a heterogeneous point cloud dataset collected by a vehicle-mounted MLS system and a helmet-mounted WLS system in Wuhan, China. High-precision vehicle-mounted MLS point clouds serve as the prior map and helmet-mounted WLS point clouds are used for localization. The dataset contains seven data: CS college, Info campus, Zhongshan park, Jiefang road 1, Jiefang road 2, Yanjiang road 1, Yanjiang road 2, as shown in Fig.7.
@@
The first two data are collected on a college campus with abundant vegetation, where roadside buildings are often obscured by trees. The road of CS college is relatively open, whereas the Info campus exhibits a more unstructured environment, with numerous blank data holes present in the prior map. The last five data are collected on urban roads with regularly arranged roadside trees, lots of high-rise buildings and viaducts in the scene, and lots of dynamic objects on the road, with variations in some of the scenes. Some of the road sections in the last five data have insufficient features, posing a significant challenge for localization. The details of creating the dataset are described in our previous work, PatchAugNet (Zou et al., 2023). It should be noted that WLS point clouds of Jiefang Road 2 is captured with Livox Mid 360, which is significantly different in scanning modes. The difference lies in the fact that the helmet point cloud submaps are generated every 0.5 m along the trajectory and the MLS point cloud submaps are derived by utilizing $5\times 5$ m regular grids. Additionally, we introduce noise into the helmet trajectory by referencing the official code of Overlap-loc (Chen et al., 2020). The comprehensive details of the experimental data are presented in Table 1.
@@
width = colspec = Q[125]Q[200]Q[81]Q[81]Q[81]Q[71]Q[71]Q[79]Q[67], cells = c, cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=20.194, cell16 = c=20.142, cell18 = r=2, cell19 = r=2, cell34 = r=7, cell35 = r=2, cell37 = r=2, cell52 = r=5, cell55 = r=5, cell89 = r=2, hline1,3,10 = -, Datasets & Scene description Difficult level Acquisition equipment Acquisition time Query
trajectory
length(km)
Number
of map
submaps
Query Map Query Map
CS college University campus road. Easy WHU-Helmet CHCNAV Alpha 3D 2022-07-29 2021-11-20 0.96 2493
Info campus University campus road, partial map is incomplete. Hard 2021-11-12 3.10 5878
Zhongshan park Urban road, the features of some road sections are insufficient. Medium Hi-Target HiScan-VUX 2022-08-23 2019-12-11 4.74 15186
Jiefang road 1 Medium 2022-08-23 2020-03-20 4.58 12202
Jiefang road 2 Hard 2024-12-13 2020-03-20 11.73 30090
Yanjiang road 1 Hard 2022-08-23 2020-03-22 2.86 13772
Yanjiang road 2 Medium 2022-08-23 2020-03-22 3.24
@@ 5.2 Experiment setting
Implementary details. We use the official model supplied by PatchAugNet (Zou et al., 2023) to extract global and local features. The spectral matching method described in Subsection 3.2 utilizes the official code of SGV (Vidanapathirana et al., 2023), renowned for its excellent parallelism capabilities. Furthermore, the local pose estimation in Subsection 4.2 is achieved based on the official code of Teaser++ (Yang et al., 2020), renowned for its robust performance. The number of particles initialized for MCL is 5000, while after convergence, it reduces to 400. In particle clustering, the distance threshold $d_{cluster}$ is set to 5 m, and $K_{thr}$ is set to 40. The parameters of the Gaussian kernel function in Eq. (11) are set to 30 m, 30 m, and 60 degrees, respectively. Low-variance resampling (Thrun, 2002) method is used for particle resampling. For localization status monitoring, the thresholds $\sigma_{x_{thr}}$, $\sigma_{y_{thr}}$, and $\sigma_{yaw_{thr}}$ are set to 30 m, 30 m, and 15 degrees, respectively. $\lambda_{thr}$ is assigned a value of 1e-4 on Yanjiang road 1, and 5e-4 for the remaining datasets. All experimental methods are implemented in Python, and all experiments are conducted on an Intel® Core i7-13700KF CPU and an Nvidia® GeForce RTX 3080Ti GPU.
Evaluation metrics. To evaluate the proposed method’s effectiveness, we compare the estimated trajectories with the ground truth trajectories and then calculate the mean error and root mean square error (RMSE) of position and yaw angle. We denote $PE$ and $YE$ as position error and yaw error respectively. It should be noted that only results after MCL converges are considered. Additionally, we count the proportion of position errors less than 2, 5, 10, 15, 20, 25, and 30 meters and yaw angle errors less than 2, 5, 10, 15, 20, 25, and 30 degrees, i.e., the localization success rate under different position and yaw error thresholds. For convenience, we denote $R@XmYd$ as the proportion of position errors less than $X$m and yaw errors less than $Y$ degrees. Similarly, $R@Xm$ denotes the proportion of position errors less than $X$m.
@@ 5.3 Quantitative evaluation
Compared with PF-loc, $R@2m5deg$, position accuracy, and yaw accuracy of PF-SGV-loc are 2.14 percent, 123.00 m, and 5.50 degrees higher, respectively. Similarly, $R@2m5deg$, position accuracy, and yaw accuracy of PF-SGV2-loc are 11.70 percent, ±108.12 m, and ±4.84 degrees higher than that of PF-SGV-loc, respectively. The reason for performance improvement is that the spatial verification-based observation model can effectively utilize the rich information embedded in local features, which is an effective complement to the observation model based on global features. Specifically, in feature-rich scenes, higher localization accuracy can be achieved with the spatial-verification-based observation model, providing a better initial pose for subsequent localization. In feature-insufficient scenes, higher localization robustness can be achieved by avoiding the particles converging to the incorrect region.
PF-SGV2-loc presents a high $R@30m$ on all data, exhibiting excellent robustness. However, the corresponding position and yaw accuracy are likely as low as ±35.15 m and ±11.65 degrees, respectively, exhibiting poor accuracy. Reg-loc achieves position and yaw accuracies of ±1.01 m and ±2.51 degrees respectively on the simple data (CS college), exhibiting excellent accuracy. While it presents a low $R@30m$ in other data with insufficient features, exhibiting poor robustness. Reliable-loc outperforms both PF-SGV2-loc and Reg-loc on all data with a total position accuracy and yaw accuracy of ±2.91 m and ±3.74 degrees, respectively, and it also presents a high $R@30m$ on all data, exhibiting excellent robustness and accuracy. The experimental results show that MCL is robust but not accurate, and localization based on point cloud registration is effective in feature-rich scenes but prone to failure in scenes with insufficient features and incomplete map coverage. However, Combining the two methods significantly improves the localization robustness and accuracy. The reasons are as follows: 1) the particles in MCL possess certain exploratory capabilities, allowing them to converge to the correct region even when the initial pose is inaccurate.; 2) After MCL converges, local pose estimation can achieve high localization accuracy in scenes with sufficient distinguishable features and provide an accurate initial pose for subsequent localization; 3) Reliable-loc can improve localization performance by adaptively switching the localization mode based on pose uncertainty from correspondences and odometry. It uses the $Reg$ localization mode when the initial pose is accurate and there are abundant features in the scene while using the $PF$ localization mode in other situations.
@@
width = colspec = Q[83]Q[100]Q[77]Q[92]Q[96]Q[92]Q[96]Q[94]Q[102]Q[96], cell21 = r=3, cell51 = r=3, cell81 = r=3, cell111 = r=3, cell141 = r=3, hline1-2,5,8,11,14,17 = -, Methods & Metrics CS college Info campus Zhongshan park Jiefang road 1 Jiefang road 2 Yanjiang road 1 Yanjiang road 2 Total
PF-loc R@2m5d (%)↑ 4.76 0.39 5.14 1.50 0.00 1.50 4.29 1.87
PE (m)↓ 8.96 ± 10.41 241.05 ± 289.59 8.28 ± 14.43 7.77 ± 10.49 556.14 ± 1628.78 7.65 ± 8.91 12.12 ± 17.06 238.90 ± 1011.32
YE (deg)↓ 7.98 ± 9.49 53.26 ± 67.94 8.25 ± 10.73 7.90 ± 10.48 17.12 ± 30.83 7.64 ± 9.24 9.90 ± 12.26 15.72 ± 28.87
PF-SGV-loc R@2m5d (%)↑ 0.81 2.33 7.00 9.24 0.00 6.58 7.96 4.01
PE (m)↓ 7.16 ± 7.95 23.49 ± 76.67 6.34 ± 12.92 5.32 ± 6.39 288.96 ± 1137.59 5.38 ± 7.27 12.66 ± 20.43 115.90 ± 701.82
YE (deg)↓ 6.68 ± 7.96 10.91 ± 22.78 7.92 ± 10.31 7.31 ± 9.34 13.15 ± 21.83 6.96 ± 8.80 9.83 ± 13.29 10.22 ± 16.94
PF-SGV2-loc R@2m5d (%)↑ 11.25 17.77 23.11 34.75 1.16 27.00 24.45 15.71
PE (m)↓ 3.38 ± 3.73 35.15 ± 102.23 3.22 ± 10.71 2.57 ± 3.16 6.98 ± 8.09 3.31 ± 4.11 4.43 ± 6.73 7.78 ± 32.37
YE (deg)↓ 4.17 ± 5.20 11.65 ± 28.36 4.38 ± 5.58 3.80 ± 4.84 5.20 ± 7.15 4.53 ± 5.97 4.91 ± 6.64 5.38 ± 10.64
Reg-loc R@2m5d (%)↑ 73.78 6.64 55.37 42.95 0.63 2.71 2.00 18.85
PE (m)↓ 1.01 ± 1.83 164.37 ± 201.16 113.08 ± 227.61 458.15 ± 773.78 594.66 ± 786.47 762.37 ± 938.22 1356.16 ± 1579.88 529.77 ± 820.08
YE (deg)↓ 2.51 ± 4.37 32.22 ± 39.78 9.82 ± 18.26 34.25 ± 54.00 33.52 ± 46.20 65.78 ± 75.52 58.02 ± 63.06 33.42 ± 47.49
@@
R@2m5d (%)↑ 77.55 52.02 80.39 72.37 30.87 35.91 68.87 53.16
PE (m)↓ 0.73 ± 1.10 2.41 ± 3.31 1.13 ± 10.29 1.92 ± 5.34 4.93 ± 6.79 2.39 ± 3.09 1.33 ± 2.59 2.91 ± 6.49
YE (deg)↓ 2.09 ± 3.55 4.02 ± 6.12 2.44 ± 4.57 2.64 ± 4.42 4.78 ± 7.05 4.45 ± 6.02 3.22 ± 5.23 3.74 ± 5.91 Note: the rightmost column ’Total’ is the overall localization performance of each method on all experimental data.
Figure 8: Localization success rate curves of each method on experimental data. (a), (b), (c), (g), (h), (i), (m) are the $R@Xm$ curves, and (d), (e), (f), (j), (k), (l), (n) are the $R@Yd$ curves.